OpenStudy (anonymous):

What are the solutions of 4x2 = -20x?

OpenStudy (anonymous):

ok so take 4x^2+20x=0, then pull out an x making it x(4x+20) and so x=0 and 4x+20=0 then x=-5 so 0 and -5